OpenStudy (anonymous):

find the function value. f(x) = -3x-4; f(0) and f(-2)

OpenStudy (anonymous):

Substitute and evaluate. For x=0\[f(x) = -3x-4\implies f(0) = -3(0)-4=-4\]You should be able to do the same thing for the other part of the problem.

OpenStudy (anonymous):

so would i go f(0) - f(-2)=2 so my answer would be -4,2

OpenStudy (anonymous):

How are you computing this? That isn't quite what I get. f(0) is the value of the function (y value) when x=0. We put zero into the function in place of x and do the arithmetic to find it. To find f(-2) (which is the y value of the function at x=-2), we do the same thing, that is, we substitute in -2 for x and do arithmetic. By my estimate\[f(x) = -3x-4\implies f(-2) = -3(-2)-4=6-4=2\]If we substitute these results into your equation that you introduced just above\[f(0) - f(-2)=-4-2=-6\]

OpenStudy (anonymous):

i don't get this at all

OpenStudy (anonymous):

I just did 0 minus -2= 2 and then I just took your -4

OpenStudy (anonymous):

I think you are sort of wandering in the dark. Do you understand the concept of function notation?

OpenStudy (anonymous):

Some good stuff about functions and function notation..... http://www.wtamu.edu/academic/anns/mps/math/mathlab/col_algebra/col_alg_tut30_function.htm